OpenStudy (anonymous):

WILL MEDAl! sqrt(5x+7) + sqrt(5x-7)

OpenStudy (anonymous):

nvm i got it 01.05 Compositions and Transformations of Functions

OpenStudy (anonymous):

cannot be simplified: \sqrt{\left(5x+7\right)}+\sqrt{\left(5x-7\right)}